# Learning Visual Composition through Improved Semantic Guidance

### v1

@@ Abstract
Visual imagery does not consist of solitary objects, but instead reflects the composition of a multitude of fluid concepts. While there have been great advances in visual representation learning, such advances have focused on building better representations for a small number of discrete objects bereft of an understanding of how these objects are interacting. One can observe this limitation in representations learned through captions or contrastive learning – where the learned model treats an image essentially as a bag of words. Several works have attempted to address this limitation through the development of bespoke learned architectures to directly address the shortcomings in compositional learning. In this work, we focus on simple, and scalable approaches. In particular, we demonstrate that by substantially improving weakly labeled data, i.e. captions, we can vastly improve the performance of standard contrastive learning approaches. Previous CLIP models achieved near chance rate on challenging tasks probing compositional learning. However, our simple approach boosts performance of CLIP substantially and surpasses all bespoke architectures. Furthermore, we showcase our results on a relatively new captioning benchmark derived from DOCCI. We demonstrate through a series of ablations that a standard CLIP model trained with enhanced data may demonstrate impressive performance on image retrieval tasks.
00footnotetext: * Equal contribution.
Correspondence: {austinstone,soltau}@google.com
@@ 1 Introduction
Figure 1: Summary of results. Previous state-of-the-art results in multimodal embeddings have limited understanding of the composition of images [23, 13] (left). The goal of this work is to learn multimodal embeddings which reflect a strong understanding of the composition of visual and semantic information (right).
@@
  • Rich visual captions are sufficient for correcting many of the failures in compositional representation (Tab. 3.1).
@@ 3.1 Beyond a bag of words representation
Table 1: Our model improves upon bag of words. Performance on Attribution, Relation, and Order (ARO) benchmark [32]. All numbers report classification macro accuracy. Chance rate is 50%. Note that BLIP and X-VLM employ a second-stage binary classification in order to exhaustively identify the best candidate. See text for details.
@@
A primary indication of the failure of multimodal representations can be observed a the relatively simple ARO evaluation dataset [32]. ARO artificially perturbs a set of captions from “the horse is eating grass” to “the grass is eating a horse” (i.e. relations) or “the paved road and the white house” to “the paved house and the white road” (i..e. attributes), and subsequently asks if the corresponding image is closer to the former or the latter caption. Chance rate is 50% and a model that does respect the relations and attributes of an image – such as a bag-of-words – would correspondingly perform at chance rate. One of the most widely used multimodal baselines, CLIP [23], achieves 59% and 63% accuracy on relations and attributes dimensions of ARO, indicating that the model is performing only slightly above a bag-of-words representation (Table 1).
@@
Admittedly, this task is artificial, and thus some captions can be correctly identified without examining an image (e.g., “grass cannot eat a horse”) merely based on the language. To test this hypothesis, we supplied Gemini 1.5 with two caption alternatives but no image. This blind baseline achieves an accuracy of 71% and 82% on relation and attributes, respectively, outperforming CLIP. This result underscores the importance of going beyond bag-of-word representations [3].
@@
Several directions have been pursued for improving performance on this baseline (Table 1). The ARO benchmark proposed a training augmentation method NegCLIP in which hard negatives are artificially added to the training set, which achieves 81% and 71% accuracy. More importantly, several works have built bespoke architectures, derived from CLIP, which exploit localization information in order to better ground visual information. Correspondingly, two different architectures achieves SOTA performance on relations (73% [33]) and attributes (88% [16]). Both models achieve these results by tieing the image and text towers together, using a grounded, cross-modal encoder [33, 16]. Although both methods employ a standard nearest neighbor lookup, they achieve SOTA results through a second-stage computation that exhaustively calculates a binary classification score for a set of candidate image-caption pairs.
@@
In comparison, our work makes no architectural changes and maintains a single contrastive training objective in a standard CLIP model. We do employ additional data augmentation with 64M synthetically generated “hard negative” examples (Section 2.1) and find that the addition of these data augmentations significantly improve ARO performance (Section 3.5). Notably, our model achieves 92% and 94% accuracy on relations and attributes surpassing bespoke architectures that exhaustively search across examples. In the Appendix, we measure how our model uniformly outperforms competing methods across the assortment of fine-grained tasks comprising the ARO benchmark.
@@ 3.3 Evaluation on detailed image caption retrieval
Given the positive results on the ARO benchmark but the inability to further discern gains on COCO, we investigated whether another retrieval benchmark may capture the gains in our approach. The DOCCI dataset was recently introduced in the context of image generation [20] but has recently been explored for image retrieval. DOCCI contains 14,847 images with detailed human-written descriptions (see [20] for details). Briefly, each caption contains on average 135.9 words across 7.1 sentences . Importantly, DOCCI purposefully contains images and corresponding captions in which there exist slight changes and alterations to the pose, orientation, background, or action. We construct an image retrieval task from this dataset and measure recall @ 1 for image retrieval. We report results on this dataset primarily in Table 2, but see Tables 4, 5, 6, and 7 for detailed ablations.
@@ 3.5 Exploring the space of improved guidance
We examine the space of semantic guidance to better understand the additive benefits of each change in our model. Given that captioning 1B images is prohibitively expensive, we perform detailed ablations on a subset of 100M images for image retrieval on DOCCI and COCO (Tables 4, 5, 6, and 7).
@@ 5 Discussion
In this work we have highlighted the failures of multimodal embeddings. Current foundational models contain limited compositional information as observed by perturbations in the attributes and relationships as well as the failure of models to perform image retrieval on challenging captioning tasks (Section 3.1, 3.2, and 3.3).
This work attempts address this problem by providing improved semantic guidance for multimodal models. In lieu of architectural innovations, this focuses on two simple and small changes that are highly scalable: (1) re-captioning training, and (2) improving the semantic target with a pretrained and powerful language model. These two improvements appear complimentary and result in a model that is able to perform 94.5 recall @1 on DOCCI image retrieval [20] (Table 2).
@@ Appendix A Prompts used for recaptioning
“The image presented came from a web page called: page title and had the alt-text: alt-text. Please very briefly describe what is in the image using the alt-text and the page title as a guide to ground your response. For example, if the alt-text contains a specific brand name, use that brand name in your output. Please describe what is in the image but be extremely concise in your response. I want to emphasize how important it is to be VERY brief. DO NOT make things up. If you can’t tell something with certainty in the image, simply don’t say anything about it.””

### v2

@@ Abstract
Visual imagery does not consist of solitary objects, but instead reflects the composition of a multitude of fluid concepts. While there have been great advances in visual representation learning, such advances have focused on building better representations for a small number of discrete objects bereft of an understanding of how these objects are interacting. One can observe this limitation in representations learned through captions or contrastive learning – where the learned model treats an image essentially as a bag of words. Several works have attempted to address this limitation through the development of bespoke architectures. In this work, we focus on simple and scalable approaches. In particular, we demonstrate that by improving weakly labeled data, i.e. captions, we can vastly improve the performance of standard contrastive learning approaches. Previous CLIP models achieved near chance rate on challenging tasks probing compositional learning. However, our simple approach boosts performance of CLIP substantially and achieves state of the art results on compositional benchmarks such as ARO and SugarCrepe. Furthermore, we showcase our results on a relatively new captioning benchmark derived from DOCCI. We demonstrate through a series of ablations that a standard CLIP model trained with enhanced data may demonstrate impressive performance on image retrieval tasks.
@@ 1 Introduction
Figure 1: Summary of results. Left: Previous state-of-the-art results in multimodal embeddings have limited understanding of the composition of images [26, 14]. Right: The goal of this work is to learn multimodal embeddings which reflect a strong understanding of the composition of visual and semantic information. Images and captions are red and blue, respectively.
@@
  • Rich visual captions are sufficient for correcting many of the failures in compositional representation (Sec. 3.1).
@@ 3.1 Beyond a bag of words representation
Table 1: Our model improves upon bag of words. Performance on Attribution, Relation, and Order (ARO) benchmark [36]. All numbers report classification macro accuracy. Chance rate is 50%. Note that BLIP and X-VLM employ a second-stage binary classification in order to exhaustively identify the best candidate. Vera and Gemini-text only see the caption and not the images.
replace swap add
O A R O A O A
Human [11] 100% 99% 97% 99% 100% 99% 99%
Vera [20] 49% 50% 49% 49% 49% 49% 50%
CLIP [13] 94% 79% 65% 60% 62% 78% 72%
DC-XL [10] 96% 85% 70% 65% 67% 91% 85%
LAION [29] 97% 86% 72% 64% 72% 93% 86%
CapPa [32] 92% 90% 87% 82% 88% 99% 99%
GPT-4V [23] 96% 94% 90% 83% 90% 92% 92%
Ours 97% 94% 88% 89% 94% 95% 93%
Table 2: Results on SugarCrepe [11] O, A, and R stand for the object, attribute, and relation split, respectively. Chance rate is 50%. Vera is a text-only model which does not see the images. GPT-4V sees both captions simultaneously and is not a two tower embedding model. LAION is xlm-roberta-large-ViT-H-14.
@@
A primary indication of the failure of multimodal representations can be observed in the relatively simple ARO [36] and SugarCrepe [11] evaluation datasets. ARO artificially perturbs a set of captions from “the horse is eating grass” to “the grass is eating a horse” (i.e. relations) or “the paved road and the white house” to “the paved house and the white road” (i..e. attributes), and subsequently asks if the corresponding image is closer to the former or the latter caption. Chance rate is 50% and a model that does respect the relations and attributes of an image – such as a bag-of-words – would correspondingly perform at chance rate. One of the most widely used multimodal baselines, CLIP [26], achieves 59% and 63% accuracy on relations and attributes dimensions of ARO, indicating that the model is performing only slightly above a bag-of-words representation (Table 1).
@@
Admittedly, this task is artificial, and some captions can be correctly identified without examining an image (e.g., “grass cannot eat a horse”) merely based on the language. To show this, we supplied Gemini 1.5 with two caption alternatives but no image. This blind baseline achieves an accuracy of 71% and 82% on relation and attributes, respectively, outperforming CLIP. SugarCrepe improves upon this deficiency such that even text-only models specifically trained for verbal plausibility such as [20] only score at random chance levels. Apart from the more plausible negative captions, SugarCrepe is similar to ARO. We additionally show our results on SugarCrepe in Table 2.
@@
Several directions have been pursued for improving performance on these baselines (Table 1, 2). The ARO benchmark proposed a training augmentation method NegCLIP in which hard negatives are artificially added to the training set, which achieves 81% and 71% accuracy. More importantly, several works have built bespoke architectures, derived from CLIP, which exploit localization information to better ground visual information. Correspondingly, two different architectures achieves SOTA performance on relations (73% [37]) and attributes (88% [17]). Both models achieve these results by tieing the image and text towers together, using a grounded, cross-modal encoder [37, 17]. Although both methods employ a standard nearest neighbor lookup, they achieve SOTA results through a second-stage computation that exhaustively calculates a binary classification score for a set of candidate image-caption pairs. [32] replaces the contrastive objective with a captioning objective and shows this leads to gains on vision & language tasks.
@@
In comparison, our work makes no architectural changes and maintains a single contrastive training objective in a standard CLIP model. We do employ additional data augmentation with 64M synthetically generated “hard negative” examples (Section 2.1) and find that the addition of these data augmentations significantly improve compositional understanding (Section 3.5). Notably, our model achieves 92% and 94% accuracy on relations and attributes surpassing bespoke architectures that exhaustively search across examples. On SugarCrepe, we achieve state of the art or highly competitive results across results across all splits, even surpassing GPT-4V [23] on all but one split. In the Appendix, we measure how our model uniformly outperforms competing methods across the assortment of fine-grained tasks comprising the ARO benchmark.
@@ 3.3 Evaluation on detailed image caption retrieval
Given the positive results on the ARO benchmark but the inability to further discern gains on COCO, we investigated whether another retrieval benchmark may capture the gains in our approach. The DOCCI dataset was recently introduced in the context of image generation [22] but has recently been explored for image retrieval. DOCCI contains 14,847 images with detailed human-written descriptions (see [22] for details). Briefly, each caption contains on average 135.9 words across 7.1 sentences . Importantly, DOCCI purposefully contains images and corresponding captions in which there exist slight changes and alterations to the pose, orientation, background, or action. We construct an image retrieval task from this dataset and measure recall @ 1 for image retrieval. We report results on this dataset primarily in Table 3, but see Tables 5, 6, 7, and 8 for detailed ablations.
@@ 3.5 Exploring the space of improved guidance
We examine the space of semantic guidance to better understand the additive benefits of each change in our model. Given that captioning 1B images is prohibitively expensive, we perform detailed ablations on a subset of 100M images for image retrieval on DOCCI and COCO (Tables 5, 6, 7, and 8).
@@ 5 Discussion
In this work we have highlighted the failures of multimodal embeddings. Current foundational models contain limited compositional information as observed by perturbations in the attributes and relationships as well as the failure to perform image retrieval on challenging captioning tasks (Section 3.1, 3.2, and 3.3).
This work addresses this problem by providing improved semantic guidance for multimodal models. In lieu of architectural innovations, this focuses on two simple and small changes that are highly scalable: (1) re-captioning training, and (2) improving the semantic target with a pretrained and powerful language model. These two improvements appear complimentary and result in a model that is able to perform 94.5 recall @1 on DOCCI image retrieval [22] (Table 3).
@@ Appendix A Prompts used for recaptioning
“The image presented came from a web page called: page title and had the alt-text: alt-text. Please very briefly describe what is in the image using the alt-text and the page title as a guide to ground your response. For example, if the alt-text contains a specific brand name, use that brand name in your output. Please describe what is in the image but be extremely concise in your response. I want to emphasize how important it is to be VERY brief. DO NOT make things up. If you can’t tell something with certainty in the image, simply don’t say anything about it."”